it is thought that a normal population's mean is less than 75. To test this, a random sample of size 19 is drawn. THe sample mean is 73.5 and the sample standard deviation is 2.1. Find an approximate range for the p-value and state the conclusion. What would your range be if we had used a two-tail test?
Answer by stanbon(75887)

it is thought that a normal population's mean is less than 75. To test this, a random sample of size 19 is drawn. THe sample mean is 73.5 and the sample standard deviation is 2.1. Find an approximate range for the p-value and state the conclusion. What would your range be if we had used a two-tail test?
------
Ho: u >= 75
Ha: u < 75 (claim)
------
Test stat: t(73.5) = (73.5-75)/[2.1/sqrt(19)] = -3.1135
-----
p(value) = P(t < -3.1135 when df = 18) = 0.0030
Since the p-value is < 5%, reject Ho at the 5% significance level.
-----
If a 2-tail test, p-value = 2*0.0030 = 0.0060
We would still reject Ho.
==================
Note: The test results support the claim.
============================================
Cheers,
Stan H.